OpenStudy (anonymous):

PLZ help what is the angle between the vectors a = (-9,8,-5) b = (2,9,-9)

OpenStudy (anonymous):

i keep getting 53.89 and that is incorrect

OpenStudy (zzr0ck3r):

a dot b = |a|*|b|cos(theta)

OpenStudy (zzr0ck3r):

so do the two vectors for the LHS then figure the magnitude of the each vector and you will solve for cos^(-1)((a dot b)/(|a||b|))

OpenStudy (anonymous):

i keep doing that and i keep getting 53.89

OpenStudy (anonymous):

i get 99/sqrt28220 every time

OpenStudy (zzr0ck3r):

and you think that is wrong?

OpenStudy (anonymous):

i know that is wrong my computer tells me so lol

OpenStudy (zzr0ck3r):

yes i get 99/(2sqrt(7055))

OpenStudy (zzr0ck3r):

sqrt(81+81+25) = sqrt(170) sqrt(4+81+81) = sqrt(166) sqrt(166)*sqrt(170) = 2sqrt(7055)

OpenStudy (zzr0ck3r):

yeah that looks liek the same thing sqrt(4*7055)

OpenStudy (anonymous):

Yeah, I'm getting 53.8907 degrees too. Does it want the answer in radians?

OpenStudy (zzr0ck3r):

does the book say 53.89?

OpenStudy (zzr0ck3r):

o hmm

OpenStudy (zzr0ck3r):

what does the book say?

OpenStudy (zzr0ck3r):

maybe it wants radians...

OpenStudy (zzr0ck3r):

is it .94 @m.auld64 ?